OpenStudy (ammarah):

The variables x and y vary directly. Write an equation that relates x and y. Then find y when x = 3.

OpenStudy (ammarah):

1. x =6, y=-8

OpenStudy (anonymous):

if y varies directly as x then y=kx where k is the constant of proportionality we need to find k first with the help of x =6, y=-8

OpenStudy (anonymous):

thus -8=k*6 or k=-4/3 hence y=(-4/3)*x

OpenStudy (anonymous):

now we need y when x=3 thus y=(-4/3)*(3)=-4

OpenStudy (ammarah):

so the answer is y=-4/3x =a and y = -4